Hi
May I know how to do the Business rate.
For 2014 were charged at a rate 0f 36000 per annum. For 2005 the rates have increased to 45000 per annum.
In both years, the annual charge is payable in 9 equal monthly instalments, from 1/4/2015
What amount is for 2014 opening and what amount is for 2015 closing.
1
Expert's answer
2020-01-24T13:20:01-0500

i=(1+0.36/9)1-1=(1+0.04)1-1=1.04-1=0.04 or 4 percent,


i=(1+0.45/9)2-1=(1+0.05)2-1=1.052-1=1.1025-1=0.1025 or 10.25 percent,


the second option is beneficial.


The interest rates are assumed in the solution.